OpenStudy (anonymous):

Evaluate. 2 – 4 • 3 + 6 • 3 A. 0 B. –28 C. 12 D. 8

OpenStudy (anonymous):

D

OpenStudy (anonymous):

A

OpenStudy (anonymous):

2-4=-2, -2x3=-6+6=0x3=0

OpenStudy (anonymous):

Hey! @mikey35633 Do you know what PEMDAS is?

OpenStudy (anonymous):

YES

OpenStudy (anonymous):

A 0

OpenStudy (anonymous):

Order of operations, its not 0. It's 8, because you have to multiply first then add